L^{A}TEX News

Issue 6, December 1996

Welcome to L^{A}TEX News 6

This issue of L^{A}TEX News accompanies the sixth release of the new standard L^{A}TEX, L^{A}TEX 2ε.

Mono-case file names

Previously L^{A}TEX has used some files with ‘mixed-case’ file names such as T1cmr.fd and T1enc.def.

These file names cause problems on some systems (in particular they are illegal on the ISO 9660 CDROM format) and so in this release all file names have been made lowercase (for example t1cmr.fd and t1enc.def).

This change should not affect any document. Within L^{A}TEX, encodings still have the usual uppercase names

in uses such as \usepackage[T1]{fontenc} and

\fontencoding{T1}. L^{A}TEX will automatically convert to the lowercase form while constructing the file name. L^{A}TEX will input the ‘fd’ file under the old name if it fails to find the file with the new name, so existing collections of fd files should still work with this new release.

The change does affect the configuration files that may be used to make the L^{A}TEX format with initex. For example, the file fonttext.ltx previously specified

\input{T1cmr.fd}. It now has \input{t1cmr.fd}. If

you use a local file fonttext.cfg you will need to make similar changes, as \input{T1cmr.fd} will not work as T1cmr.fd is no longer in the distribution.

The files affected by this change all have names of the

form *.fd or *enc.def.

Another input encoding

Thanks to work by Søren Sandmann, the inputenc package now supports the IBM codepage 865 used in Scandinavia.

Better user-defined math display environments

Suppose that you want to define an environment for displaying text that is numbered as an equation. A straightforward way to do this is as follows:

  \newenvironment{texteqn}
    {\begin{equation}
       \begin{minipage}{0.9\linewidth}}
      {\end{minipage}
     \end{equation}}

However, if you have tried this then you will probably have noticed that it does not work perfectly when used in the middle of a paragraph because an inter-word space appears at the beginning of the first line after the environment.

There is now an extra command (with a very long name) available that you can use to avoid this problem; it should be inserted as shown here:

                                                     \newenvironment{texteqn}
                                                       {\begin{equation}
                                                          \begin{minipage}{0.9\linewidth}}
                                                         {\end{minipage}
                                                        \end{equation}
                                                        \ignorespacesafterend}

Docstrip improvements

The docstrip program that is used to unpack the L^{A}TEX sources has undergone further development. The new version should be able to process all old ‘batchfiles’ but it allows a simpler syntax in new ‘batchfiles’ (no need to

                                                   define \def\batchfile{. . . ).

It also allows ‘target’ directories to be specified when writing files. This directory support is disabled by default unless activated in a local docstrip.cfg configuration file. See docstrip.dtx for details.

AMS L^{A}TEX update

Since the last L^{A}TEX release in June, the American Mathematical Society have re-issued the ‘AMSL^{A}TEX’ classes and packages, fixing several reported problems.

Graphics package update

The L^{A}TEX color and graphics packages have been updated slightly, principally to support more dvi drivers, see the readme file in the graphics distribution.

EC Fonts released

The first release of the Extended Computer Modern fonts has just been made. (In January 1997.)

This release of L^{A}TEX does not default to these ‘ec’ fonts as its T1 encoded fonts. By default it will use the ‘dc’ fonts if the T1 encoding is requested.

As noted in install.txt you may run TEX on the install file ec.ins after unpacking the base distribution but before making the L^{A}TEX format. This will produce suitable ‘fd’ files making L^{A}TEX (including, for the first time, the slides class) use the ‘ec’ fonts as the default T1 encoded font set.

L^{A}TEX News, and the L^{A}TEX software, are brought to you by the L^{A}TEX3 Project Team; Copyright 1996, all rights reserved.